OpenStudy (anonymous):

The shorter leg of a 30°- 60°- 90° right triangle is 12.5 inches. How long are the longer leg and the hypotenuse?

OpenStudy (anonymous):

ratios of a 30 - 60 - 90 right triangle are \(1:\sqrt{3}:2\)

OpenStudy (anonymous):

so if the short let is say \(a\) then the long leg is \(a\times \sqrt{3}\) and the hypotenuse is \(2a\)

OpenStudy (anonymous):

so the longer leg would be 21.65 and the hypotenuse would be 25

OpenStudy (raden):

that's right ^

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

yw